ULTRAHIGH-RESOLUTION SPECTROSCOPY OF DIBENZOFURAN S_{1} ← S_{0} TRANSITION

\underline{SHUNJI KASAHARA} and MICHIRU YAMAWAKI, Molecular Photoscience Research Center, Kobe University, Kobe 657-8501, Japan; MASAAKI BABA, Graduate School of Science, Kyoto University, Kyoto 606-8502, Japan.

Dibenzofuran is one of the prototypical molecule of toxic dioxins and it is great important to investigate the excited-state dynamics. We have observed low-resolution spectra for the vibronic structure of the S_{1} ^{1}A_{1} ← S_{0} ^{1}A_{1} transition of jet-cooled dibenzofuran. Several vibronic bands are stronger than the 0_{0}^{0} band. These bands are found to be the A-type transition which are enhanced by intensity borrowing through the vibronic coupling with the S_{2} ^{1}B_{2} state. ^{a} Rotationally resolved ultrahigh-resolution spectra of jet-cooled dibenzofuran for the 0_{0}^{0} and 0_{0}^{0}+443 cm^{−1} band have been observed. Several thousand lines were assigned and these molecular constants were determined. ^{b} We have extended the highresolution spectroscopic measurement for several vibronic bands up to 1335 cm^{−1}. We analyzed observed rotational lines and determined the molecular constants for each vibronic bands. The Zeeman broadenings of the rotational lines were also observed, and the J, K-dependence were studied. As a result, the observed Zeeman effects could be explained to be originating from the magnetic moment of the S_{1} ^{1}A_{1} state induced by mixing with S_{2} ^{1}B_{2} state by J-L coupling.

aM. Baba, K. Mori, M. Yamawaki, K. Akita, M. Ito, S. Kasahara, and T. Yamanaka, J. Phys. Chem. A 110, 10000 (2006).

bM. Yamawaki, Y. Tatamitani, A. Doi, S. Kasahara, and M. Baba, J. Mol. Spectrosc., 238, 49 (2006).